# When Federated Learning Meets Quantum Computing: Survey and Research Opportunities

### v2

@@ 4.2.1 Models with Architectural Integration
Architectural integration forms the backbone of QFL, defining how quantum and classical components collaborate to achieve distributed learning objectives. Hybrid architectures [55], such as those that utilize VQCs or tensor-based networks, emphasize communication efficiency and scalability. These models often achieve significant reductions in communication overhead while maintaining high performance on specific tasks [56]. However, challenges persist in generalizing these architectures to heterogeneous and dynamic environments, particularly when addressing the trade-offs between communication optimization and model accuracy. Previous works highlight the need for robust architectural frameworks that balance computational demands with adaptability to real-world conditions [57].
@@ 4.2.2 FL on NISQ Devices
The integration of FL on NISQ devices represents a critical dimension, focusing on the practical deployment of FL within the constraints of quantum hardware. Models designed for NISQ devices often prioritize noise resilience and resource optimization, with techniques such as shallow circuits and error mitigation strategies playing a pivotal role. While prior surveys [24, 25, 26, 27, 28, 29, 30, 31, 29, 32] have elaborated on performance metrics, including accuracy rates in noise-prone environments, the scalability of such systems remains a pressing challenge. This category analyzes the importance of developing hardware-aware algorithms that can accommodate the variability and limitations of NISQ devices, paving the way for scalable QFL.
@@ 4.2.3 Privacy-Preserving QFL
Privacy preservation is a cornerstone of QFL, addressing the dual objectives of securing sensitive data and ensuring robust federated training. Techniques such as quantum-enhanced differential privacy and homomorphic encryption have emerged as promising solutions [58, 59], achieving strong privacy guarantees while maintaining competitive model performance. However, the computational overhead and trade-offs associated with these methods, particularly in noisy quantum environments, pose significant challenges. This dimension highlights the need for lightweight and scalable privacy-preserving mechanisms that can be seamlessly integrated into FL frameworks, ensuring security without compromising efficiency.
@@ 5 Models with Architectural Integration
In QFL, models with architectural integration [60, 61, 62, 63, 64, 54, 52, 49]represent a critical nexus between QC and distributed ML. These models aim to synergize quantum-enhanced computation with classical frameworks, promising unprecedented gains. The crux lies in integration, where the models leverage quantum advantages without falling prey to hardware limitations. This demands a delicate balance, optimizing theoretical quantum benefits against practical constraints while maintaining the core principles of data sovereignty and collaborative learning. Some common and useful terms are defined below:
@@
$\bullet$ Challenges: QFL models offer theoretical advantages, but architectural integration is challenging. A major issue is noise and decoherence in quantum devices, which reduces computational accuracy and stability [65]. VQCs, key to hybrid models, often face barren plateaus, limiting scalability in deep quantum networks [66]. The probabilistic nature of quantum measurements adds variability to gradient updates, complicating optimization and convergence in federated training [52]. These challenges hinder the alignment of the quantum and classical frameworks while maintaining federated privacy [54].
Communication bottlenecks, especially in federated quantum settings where resources are scarce, hinder progress. Quantum-assisted models require frequent parameter exchanges through noisy channels, increasing quantum state transmission costs [67]. Integrating quantum tensor decompositions and hybrid layer fusion demands optimized circuit depths to avoid decoherence and high resource use [68]. While tensor-network circuits and hierarchical updates can reduce complexity, their use is limited by the shortage of fault-tolerant qubits [69].
@@ 5.1 Quantum-Classical Layer Fusion
Quantum-classical layer fusion [70, 71] constitutes a fundamental architectural paradigm in hybrid learning models. This approach ingeniously replaces specific neural network layers, particularly those demanding substantial computational resources, with quantum equivalents. The quantum layers exploit the unique phenomena of quantum mechanics, superposition and entanglement, to execute more intricate transformations on the input data. Concurrently, classical layers manage less resource-intensive operations, ensuring the model’s viability for real-world applications. This synergistic integration of quantum and classical elements yields a hybrid neural network that, in theory, accelerates learning and enhances model accuracy [55]. QNNs (shown in Fig. 6) are often employed in such architectures, leading to lightweight FL models [72]. This fusion architecture thus represents a critical step towards harnessing quantum computational advantages while navigating the constraints of current quantum hardware, potentially revolutionizing FL across diverse domains.
@@
[49] Effective use of QCNNs for scalable learning; first implementation of quantum weights High noise sensitivity and limited real-world privacy-preservation High [shallow circuit design] Asynchronous aggregation
@@
[50] Pure QFL framework with quantum data, utilizing QCNN Potential scalability issues; vulnerable to attacks Moderate [8-qubit QCNN circuits] Asynchronous aggregation
@@
[52] High efficiency in training and reduced communication rounds Poor robustness with non-uniform and complex data distributions High [variational circuit design] Synchronous aggregation
@@
[54] Comprehensive comparative analysis across FL frameworks; includes quantum network security insights Lacks practical benchmarks or real deployment results Moderate [PQC-based circuits] Hybrid aggregation
@@
[64] Strong privacy mechanisms for non-IID data; adaptable to distributed systems High communication overhead and potential scalability issues High [efficient VQA circuits] Asynchronous aggregation
@@
[63] Efficient handling of communication and computational cost Limited adaptability in highly distributed or large-scale settings Moderate [tensor decomposition strategies] Synchronous aggregation
@@
[62] High data privacy integration; effective federated updates Complex implementation may limit practical scalability Moderate [hierarchical tensor strategies] Hybrid aggregation
@@
[61] Efficient for time-series data; reduced training overhead Low adaptability to non-time-series data and basic noise handling High [optimized QLSTM structure] Asynchronous aggregation
@@
[60] Robust noise resilience; adaptive training using fuzzy logic High complexity in setup and limited broad dataset testing Moderate [adaptive QFNNs] Asynchronous aggregation
@@
One key advantage of layer fusion is that it does not require fully developed QNNs. This is beneficial as this fusion may prove to be infeasible due to the limitations of existing quantum hardware. Instead, by partitioning the neural network, only specific layers leverage quantum resources, making this a more practical and scalable solution. An alternative solution involves the implementation of a weighted personalization algorithm that addresses data heterogeneity and model drift [73]. Recent studies demonstrated that hybrid models that use this technique can outperform fully classical models in specific tasks such as quantum chemistry simulations and optimization problems [74, 75]. Despite its promise, this model is still heavily constrained by the need for efficient quantum gates and error-correcting mechanisms that are not yet mature [76].
@@
An examination of the QUE values in Table III shows that most layer-fused models operate with moderate to high qubit‐layer occupancy (typically in the range of 67% to 100%). Shallow QCNN hybrids like [49] and [50] exploit asynchronous updates to capitalize on their high QUE, since clients can independently push low-latency gradient estimates without waiting for global synchronization. In contrast, approaches with deeper variational circuits, such as [52] and [63], favor synchronous aggregation to ensure that their more complex parameter landscapes converge cohesively. Despite these strengths, layer fusion introduces new challenges. The heterogeneity of classical and quantum execution pipelines can cause unpredictable latency spikes, complicating the coordination of model updates. Furthermore, the lack of standardized frameworks for hybrid network deployment often leads to integration overhead.
@@ 5.2 Quantum-Assisted Transfer Learning
Quantum-assisted transfer learning [49, 77] introduces the potential to enhance classical Transfer Learning (TL) techniques by leveraging quantum computing for feature extraction and optimization. This approach enables better generalization across tasks using quantum algorithms to identify and transfer rich feature representations.
A notable research in [49] explored Federated Quantum ML, utilizing hybrid quantum-classical models. By combining quantum circuits with pre-trained classical models, this work accelerates convergence without sacrificing performance, significantly addressing the scalability challenge in NISQ devices [23]. The proposed methodology is also the first to implement quantum weights, enabling a fully quantum ring network topology for QFL. Consequently, this network structure eliminates the need for quantum memory, supporting the no-cloning theorem. However, a key challenge lies in balancing communication overhead between quantum nodes in a federated system, especially when dealing with quantum noise and error propagation across the network. This model is also open to membership inference and related attacks, which raises concerns about its privacy preservation. Another work [77] proposed a novel TL architecture, which leverages quantum networks and pre-trained classical deep neural networks, to improve the adaptation of the model between related tasks. The architecture further uses VQCs for encoding and transforming classical data into quantum representations, thereby facilitating faster knowledge transfer. A critical insight from this study is that quantum states can enhance the capacity for information transfer, particularly when dealing with high-dimensional data. However, the inherent noise in quantum systems and the limitations of quantum hardware pose challenges to stable model performance during transfer learning. Furthermore, this model also suffers from privacy leakage during the exchange of model parameters, which can be used to replicate the training data.
An assessment of $\mathrm{QUE}$ for the transfer‐learning schemes suggests moderate to high utilization of qubit–layer resources. In [49], shallow QCNN circuits (e.g. $\mathrm{Q}_{alloc}=4$, $L=4$) yield $\mathrm{QUE}\approx 1.0$, which pairs well with its asynchronous update strategy to minimize idle time on quantum nodes. By contrast, the hybrid ring architecture in [77] employs slightly deeper VQCs (e.g. $\mathrm{Q}_{alloc}=4$, $L=5$), giving $\mathrm{QUE}\approx 0.80$; this design favors synchronous aggregation so that rich quantum features are consistently merged with classical weights each round. Beyond general communication and privacy concerns, TL introduces subtle model‐drift risks when quantum‐enhanced representations misalign with downstream classical layers. Moreover, the dynamic adaptation of pre-trained weights can amplify decoherence effects if quantum layers are not re-tuned post-transfer, underscoring the need for co-optimization of classical and quantum learning rates.
@@ 5.3 Federated Quantum Embedding
Federated Quantum Embedding (FQE) [50, 54] represents a cutting-edge approach in which classical data are encoded into quantum states, allowing richer and more expressive feature representations across FL systems. This method enables quantum devices to extract complex features from classical datasets by leveraging the high-dimensionality of quantum states, which can also be shared and aggregated in an FL framework.
@@
The work in [50] proposed a framework for processing purely quantum data in QFL, demonstrating that quantum embeddings can significantly reduce the communication cost associated with distributed learning. The key advantage of FQE lies in its ability to compress high-dimensional classical data into compact quantum states by employing techniques such as quantum tensor train decomposition [78], thus optimizing both communication efficiency and learning performance. The proposed model implements the Quantum Convolutional Neural Networks (QCNN)-QFL framework using the first federated quantum dataset. However, the challenge remains to manage the quantum noise introduced during the embedding process, which can degrade the quality of the shared quantum states, thus necessitating robust error mitigation techniques to ensure stable model updates. Furthermore, testing the model on larger and more practical datasets is an open challenge. Building on this, another study [54] investigated FQE in classical and quantum networks, emphasizing its potential to boost learning in federated settings with large-scale distributed data. It examines efficient quantum embedding aggregation and tackles coherence loss during data transmission. Furthermore, novel quantum communication protocols are suggested to maintain data integrity in noisy channels, offering a promising solution to the challenge of noisy quantum channels [79]. By encoding sensitive data into quantum states, FQE is inherently harder for adversaries to reverse-engineer the original information from the embeddings, thus enhancing the privacy guarantees of FL systems. However, ongoing work is needed to address decoherence and develop scalable error correction to preserve quantum embedding fidelity over long communication distances. Moreover, the lack of empirical results leaves scope for further experimentation.
@@
An evaluation of $\mathrm{QUE}$ for the embedding‐based models shows that the QCNN‐QFL ($\mathrm{Q}_{alloc}=4$, $L=3$) framework of [50] achieves $\mathrm{QUE}=0.80$, reflecting its compact three‐layer encoding circuit and justifying its asynchronous update strategy, which minimizes idle qubit time during on‐demand embedding. In contrast, the comparative study of [54] across classical and quantum networks uses a slightly deeper ansatz ($\mathrm{Q}_{alloc}=6$, $L=4$), yielding $\mathrm{QUE}=0.75$, and favors a hybrid classical–quantum fusion approach to merge classical gradient averages with quantum embedding deltas. Apart from embedding noise and coherence loss, these schemes confront the overhead of repeated state preparation and measurement for each data batch, which can dominate communication cost as dataset sizes grow. Furthermore, the dependence on accurate quantum‐classical interface calibration introduces a brittle coupling between the encoding circuit and downstream classical layers, suggesting a need for adaptive calibration protocols and lightweight embedding circuits.
@@
  • Through TL, quantum circuits can be used for feature extraction and improving model generalization in FL tasks. This approach can speed up convergence and enhance performance on non-IID datasets, but is limited by communication overhead and may not be noise resilient.

### v4

@@ 4.2.1 Models with Architectural Integration
Architectural integration forms the backbone of QFL, defining how quantum and classical components collaborate to achieve distributed learning objectives. Hybrid architectures [Fusion1], such as those that utilize VQCs or tensor-based networks, emphasize communication efficiency and scalability. These models often achieve significant reductions in communication overhead while maintaining high performance on specific tasks [ferrari2021compiler]. However, challenges persist in generalizing these architectures to heterogeneous and dynamic environments, particularly when addressing the trade-offs between communication optimization and model accuracy. Previous works highlight the need for robust architectural frameworks that balance computational demands with adaptability to real-world conditions [hisamori2024hybrid].
@@ 4.2.2 FL on NISQ Devices
The integration of FL on NISQ devices represents a critical dimension, focusing on the practical deployment of FL within the constraints of quantum hardware. Models designed for NISQ devices often prioritize noise resilience and resource optimization, with techniques such as shallow circuits and error mitigation strategies playing a pivotal role. While prior surveys [narottama2020quantum, huang2022quantum, ren2023towards, gurung2023quantum, chehimi2024foundations, quy2024from, QFL-type, javeed2023quantum, quy2024from, saha2024multifaceted] have elaborated on performance metrics, including accuracy rates in noise-prone environments, the scalability of such systems remains a pressing challenge. For many quantum systems, additional qubits are used for error correction. While early-stage techniques like repetition codes and surface codes can partially reduce noise in NISQ systems, their scaling can present significant challenges in FL environments, which are detailed in Section 6. This category analyzes the importance of developing hardware-aware algorithms that can accommodate the variability and limitations of NISQ devices.
@@ 4.2.3 Privacy-Preserving QFL
Privacy preservation is a cornerstone of QFL, addressing the dual objectives of securing sensitive data and ensuring robust federated training. Techniques such as quantum-enhanced differential privacy and homomorphic encryption have emerged as promising solutions [ullah2024quantum, broadbent2015quantum], achieving strong privacy guarantees while maintaining competitive model performance. However, the computational overhead and trade-offs associated with these methods, particularly in noisy quantum environments, pose significant challenges.
@@ 5 Models with Architectural Integration
In QFL, models with architectural integration [QFL-QFuzzyNN, QFL-QLSTM, QFL-Tensor, QFL-Tensor2, QFL-QAOA, FQE2, QFL-QNGD, QFL-TL] represent a critical nexus between QC and distributed ML. These models aim to synergize quantum-enhanced computation with classical frameworks, promising unprecedented gains. The crux lies in integration, where the models leverage quantum advantages without falling prey to hardware limitations. This demands a delicate balance, optimizing theoretical quantum benefits against practical constraints while maintaining the core principles of data sovereignty and collaborative learning. Some common and useful terms are defined below:
@@
$\bullet$ Challenges: QFL models offer theoretical advantages, but architectural integration is challenging. A major issue is noise and decoherence in quantum devices, which reduces computational accuracy and stability [shnirman2002noise]. VQCs, key to hybrid models, often face barren plateaus, limiting scalability in deep quantum networks [mcclean2018barren]. The probabilistic nature of quantum measurements adds variability to gradient updates, complicating optimization and convergence in federated training [QFL-QNGD]. These challenges hinder the alignment of the quantum and classical frameworks [FQE2].
Communication bottlenecks, especially in federated quantum settings where resources are scarce, hinder progress. Quantum-assisted models require frequent parameter exchanges through noisy channels, increasing quantum state transmission costs [streltsov2012quantum]. Integrating quantum tensor decompositions and hybrid layer fusion demands optimized circuit depths to avoid decoherence and high resource use [gelss2022low]. While tensor-network circuits and hierarchical updates can reduce complexity, their use is limited by the shortage of fault-tolerant qubits [anshu2024circuit].
@@ 5.1 Quantum-Classical Layer Fusion
Quantum-classical layer fusion [chen2024research, domingo2024hybrid] constitutes a fundamental architectural paradigm in hybrid learning models. This approach ingeniously replaces specific neural network layers, particularly those demanding substantial computational resources, with quantum equivalents. The quantum layers exploit the unique phenomena of quantum mechanics, superposition, and entanglement, to execute more intricate transformations on the input data. Concurrently, classical layers manage less resource-intensive operations, ensuring the model’s viability for real-world applications. This synergistic integration of quantum and classical elements yields a hybrid neural network that, in theory, accelerates learning and enhances model accuracy [Fusion1]. QNNs (shown in Fig. 6) are often employed in such architectures, leading to lightweight FL models [park2025quantum]. This fusion architecture thus represents a critical step towards harnessing quantum computational advantages while navigating the constraints of current quantum hardware.
@@
However, in practical deployments, fusion is not a trivial layer substitution but rather the careful placement of quantum modules at strategic points in the network. Each fusion point dictates the volume of quantum-classical communication, which, in a federated setting, directly affects aggregation latency and susceptibility to decoherence during data exchange. NISQ-era limitations remain a decisive factor: a restricted qubit allocation $Q_{\text{alloc}}$ constrains expressivity, deeper circuits amplify decoherence, and gate infidelity injects noise into the gradient update process. Moreover, bidirectional transfers between quantum processors and classical nodes increase round-trip latency.
Potential applications: Fusion-based architectures have already been adopted in several domains. In wearable healthcare analytics, quantum recurrent layers such as quantum LSTM have been integrated to improve energy expenditure estimation under constrained IoT device resources [tran2025quantumlstm]. In integrated satellite–ground remote sensing, quantum-enhanced convolutional layers can distill multi-modal features from heterogeneous sensors, thereby reducing the classical bandwidth requirement for model synchronization [khalid2025quantumfusion]. Privacy-sensitive NLP tasks can benefit from quantum embedding layers fused with classical attention mechanisms, enabling richer contextual encodings that are inherently more resistant to gradient inversion attacks.
[…]
$\mathbf{h}_{q}=\langle 0^{\otimes Q_{\text{alloc}}}|U^{\dagger}(\boldsymbol{\theta})\,\hat{O}\,U(\boldsymbol{\theta})|0^{\otimes Q_{\text{alloc}}}\rangle,$} (6)
where $\hat{O}$ is the measurement observable. The resulting $\mathbf{h}_{q}$ is concatenated with classical activations and propagated through subsequent layers in the FL pipeline, ensuring that quantum-derived features influence the global aggregation. Fig. 7 illustrates this architecture, where PQCs replace targeted classical layers and exchange measurement statistics with the classical stack during training.
Figure 7: Illustrating quantum and classical layer fusion with its limitations (red-colored) and advantages (green-colored).
@@
[QFL-TL] Effective use of QCNNs for scalable learning; first implementation of quantum weights High noise sensitivity and limited real-world privacy-preservation High [shallow circuit design] Asynchronous aggregation
@@
[FQE1] Pure QFL framework with quantum data, utilizing QCNN Potential scalability issues; vulnerable to attacks Moderate [8-qubit QCNN circuits] Asynchronous aggregation
@@
[QFL-QNGD] High efficiency in training and reduced communication rounds Poor robustness with non-uniform and complex data distributions High [variational circuit design] Synchronous aggregation
@@
[FQE2] Comprehensive comparative analysis across FL frameworks; includes quantum network security insights Lacks practical benchmarks or real deployment results Moderate [PQC-based circuits] Hybrid aggregation
@@
[QFL-QAOA] Strong privacy mechanisms for non-IID data; adaptable to distributed systems High communication overhead and potential scalability issues High [efficient VQA circuits] Asynchronous aggregation
@@
[QFL-Tensor2] Efficient handling of communication and computational cost Limited adaptability in highly distributed or large-scale settings Moderate [tensor decomposition strategies] Synchronous aggregation
@@
[QFL-Tensor] High data privacy integration; effective federated updates Complex implementation may limit practical scalability Moderate [hierarchical tensor strategies] Hybrid aggregation
@@
[QFL-QLSTM] Efficient for time-series data; reduced training overhead Low adaptability to non-time-series data and basic noise handling High [optimized QLSTM structure] Asynchronous aggregation
@@
[QFL-QFuzzyNN] Robust noise resilience; adaptive training using fuzzy logic High complexity in setup and limited broad dataset testing Moderate [adaptive QFNNs] Asynchronous aggregation
@@
One key advantage of layer fusion is that it does not require fully developed QNNs. This is beneficial as this fusion may prove to be infeasible due to the limitations of existing quantum hardware. Instead, by partitioning the neural network, only specific layers leverage quantum resources, making this a more practical and scalable solution. An alternative solution involves the implementation of a weighted personalization algorithm that addresses data heterogeneity and model drift [gurung2025performance]. Recent studies demonstrated that hybrid models that use this technique can outperform fully classical models in specific tasks such as quantum chemistry simulations and optimization problems [Chem1, optimization1]. Despite its promise, this model is still heavily constrained by the need for efficient quantum gates and error-correcting mechanisms that are not yet mature [error1].
@@
An examination of the QUE values in Table III shows that most layer-fused models operate with moderate to high qubit‐layer occupancy (typically in the range of 67% to 100%). Shallow QCNN hybrids like [QFL-TL] and [FQE1] exploit asynchronous updates to capitalize on their high QUE, since clients can independently push low-latency gradient estimates without waiting for global synchronization. In contrast, approaches with deeper variational circuits, such as [QFL-QNGD] and [QFL-Tensor2], favor synchronous aggregation to ensure that their more complex parameter landscapes converge cohesively. Despite these strengths, layer fusion introduces new challenges. The heterogeneity of classical and quantum execution pipelines can cause unpredictable latency spikes.
@@ 5.2 Quantum-Assisted Transfer Learning
Quantum-assisted transfer learning [QFL-TL, QFL-TL-2] introduces the potential to enhance classical Transfer Learning (TL) techniques by leveraging quantum computing for feature extraction and optimization. This approach enables better generalization across tasks using quantum algorithms to identify and transfer rich feature representations.
A notable research in [QFL-TL] explored Federated Quantum ML, utilizing hybrid quantum-classical models. By combining quantum circuits with pre-trained classical models, this work accelerates convergence without sacrificing performance, significantly addressing the scalability challenge in NISQ devices [Preskill2018]. The proposed methodology is also the first to implement quantum weights, enabling a fully quantum ring network topology for QFL. Another work [QFL-TL-2] proposed a novel TL architecture, which leverages quantum networks and pre-trained classical deep neural networks, to improve the adaptation of the model between related tasks. A critical insight from this study is that quantum states can enhance the capacity for information transfer, particularly when dealing with high-dimensional data.
An assessment of $\mathrm{QUE}$ for the transfer‐learning schemes suggests moderate to high utilization of qubit–layer resources. In [QFL-TL], shallow QCNN circuits (e.g. $\mathrm{Q}_{alloc}=4$, $L=4$) yield $\mathrm{QUE}\approx 1.0$, which pairs well with its asynchronous update strategy to minimize idle time on quantum nodes. By contrast, the hybrid ring architecture in [QFL-TL-2] employs slightly deeper VQCs (e.g. $\mathrm{Q}_{alloc}=4$, $L=5$), giving $\mathrm{QUE}\approx 0.80$; this design favors synchronous aggregation so that rich quantum features are consistently merged with classical weights in each round.
@@ 5.3 Federated Quantum Embedding
Federated Quantum Embedding (FQE) [FQE1, FQE2] represents a cutting-edge approach in which classical data are encoded into quantum states, allowing richer and more expressive feature representations across FL systems. This method enables quantum devices to extract complex features from classical datasets by leveraging the quantum states, which can also be shared and aggregated in an FL framework.
@@
The work in [FQE1] proposed a framework for processing purely quantum data in QFL, demonstrating that quantum embeddings can significantly reduce the communication cost associated with distributed learning. The key advantage of FQE lies in its ability to compress high-dimensional classical data into compact quantum states by employing techniques such as quantum tensor train decomposition [bhatia2025application], thus optimizing both communication efficiency and learning performance. The proposed model implements the Quantum Convolutional Neural Networks (QCNN)-QFL framework using the first federated quantum dataset. However, the challenge remains to manage the quantum noise introduced during the embedding process, which can degrade the quality of the shared quantum states, thus necessitating robust error mitigation techniques to ensure stable model updates. Building on this, another study [FQE2] investigated FQE in classical and quantum networks, emphasizing its potential to boost learning in federated settings with large-scale distributed data. It examines efficient quantum embedding aggregation and tackles coherence loss during data transmission. Furthermore, novel quantum communication protocols are suggested to maintain data integrity in noisy channels, offering a promising solution to the challenge of noisy quantum channels [chu2023cryptoqfl]. By encoding sensitive data into quantum states, FQE is inherently harder for adversaries to reverse-engineer the original information from the embeddings, thus enhancing the privacy guarantees.
@@
An evaluation of $\mathrm{QUE}$ for the embedding‐based models shows that the QCNN‐QFL ($\mathrm{Q}_{alloc}=4$, $L=3$) framework of [FQE1] achieves $\mathrm{QUE}=0.80$, reflecting its compact three‐layer encoding circuit and justifying its asynchronous update strategy, which minimizes idle qubit time during on‐demand embedding. In contrast, the comparative study of [FQE2] across classical and quantum networks uses a slightly deeper ansatz ($\mathrm{Q}_{alloc}=6$, $L=4$), yielding $\mathrm{QUE}=0.75$, and favors a hybrid classical–quantum fusion approach to merge classical gradient averages with quantum embedding deltas. Apart from embedding noise and coherence loss, these schemes confront the overhead of repeated state preparation and measurement for each data batch, which can dominate communication cost as dataset sizes grow.
@@
  • Through TL, quantum circuits can be used for feature extraction and improving model generalization in FL tasks. This approach can speed up convergence and enhance performance on non-IID datasets, but is limited by communication overhead.